### v2

# Deep Neural OFDM Receivers: Two Novel Architectures for BER-BLER Optimization and Comparison with State-of-the-Art Architectures

Abstract
Neural receivers have recently become a popular topic, where the received signals can be directly decoded by data-driven mechanisms such as machine learning and deep learning. In this paper, we propose two novel neural network–based orthogonal frequency division multiplexing (OFDM) receivers performing channel estimation and equalization tasks and directly predicting log-likelihood ratios (LLRs) from the received in-phase and quadrature-phase (IQ) signals. The first network, the Dual Attention Transformer (DAT), employs a state-of-the-art (SOTA) transformer architecture with an attention mechanism. The second network, the Residual Dual Non-Local Attention Network (RDNLA), utilizes a parallel residual architecture with a non-local attention block. The bit-error rate (BER) and block-error rate (BLER) performance of various SOTA neural receiver architectures is compared with our proposed methods across different signal-to-noise ratio (SNR) levels. The simulation results show that DAT and RDNLA outperform both traditional communication systems and existing neural receiver models. The computational efficiency of the proposed neural receivers supports their feasibility for next-generation communication systems.
Index Terms:
Neural Receivers, LLR, Transformers
Data-driven techniques such as machine learning (ML) and deep learning (DL) have seen significant advances recently, not only in foundational algorithms but also in hardware design perspective, mainly driven by the “artificial intelligence (AI) at the edge” concept. For example, tensor accelerators along with new AI algorithms make computationally challenging tasks feasible to be deployed in real-world applications. These advancements facilitate an increasing number of AI deployements at the edge [1]. Wireless communication community has also been heavily investing AI/ML based applications, where they can either build new foundational models or adopt existing ones from other domains by training them using wireless communication datasets.
Considering advanced radio access network (RAN) technologies along with the usage of larger bandwidths in the upcoming 3GPP releases, signal processing requirements of enormous data in real-time make a traditional mathematical-based system design with cascaded communication blocks consisting heuristic algorithms prohibitively expensive. This brings an AI-native system design forefront, where the AI/ML approaches can focus on the most important features of the underlying system by appropriately setting the weights of the trained models which provide a direct mapping between transmitted (input) and received (output) information. In addition, the AI/ML models are suitable for revealing highly complex features for several advanced RAN functions, where hardware accelerators play a critical role in processing enormous data in parallel [2]. As long as high-quality data is available, the model is trained and deployed at the edge. The 3GPP has already started to study AI/ML adaptation for wireless systems [3].
Neural receivers have been received significant attention recently, where signals are decoded by performing channel estimation, equalization, and symbol de-mapping tasks of the conventional receivers by the AI model. For example, in [4], a fully convolutional deep neural network, DeepRx, is introduced to execute the whole receiver pipeline from frequency domain signal processing to channel decoding. Similarly, in [5], HybridDeepRx proposes residual deep learning modules including both time and frequency domain operations for demodulating orthogonal frequency division multiplexing (OFDM) signals under challenging conditions such as a high level of non-linear distortion. Furthermore, it is shown in [6] that an end-to-end neural receiver for OFDM systems under realistic fading channels can significantly reduce pilot overhead by providing a similar bit-error rate (BER) performance compared to conventional pilot-based systems.
The recent studies on a tensor native wireless communication [7] and ray tracing based simulation framework Sionna [8] provide important advancements in the context of neural receivers. The potential of neural receiver architectures has been demonstrated in [9] and [10], where neural receivers are utilized in 5G multi-user multiple input-multiple output (MIMO) scenarios in real-time while adhering to the standard compliance. Deep learning-based channel estimation methods [11], incorporating belief knowledge and transfer learning strategies [12] for single input-multiple output (SIMO) receivers [13] further improve performance by adapting to realistic channel conditions, as verified by aerial experiments [13]. In addition, interpretability analysis [14] and performance verification frameworks for AI-specific transceiver actions [15] provide important guidance for the robustness, safety, and operational reliability of these models. Also, [16] showed neural receiver performance can be further improved by adopting new DL models such as transformers.
Motivated by these, we propose two novel neural receiver architectures developed in Python using the Sionna framework that can be adapted to standard-based wireless communication systems. The main contributions of the paper can be summarized as:
Figure 1: System Model of conventional and neural network based receiver.
  • We introduce two deep learning models for neural receivers that significantly improve BER and block-error rate (BLER) performance compared to the conventional Transformers. Proposed models also demonstrate notable improvements over other state-of-the-art (SOTA) models.
  • We then explore the architecture of these models, carefully going over each element of their construction. We highlight how each element contributes to feature extraction and learning at the global and local levels that eventually affect BER and BLER.
  • Lastly, we will make publicly available the source code of our proposed models together with other SOTA model implementations [^1] https://github.com/Hisar-Research/sionna-nn-receiver-playground. This allows researchers and practitioners to explore, test, and expand upon our findings.
Notation: Throughout this paper, vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), and matrices or tensors by bold uppercase letters (e.g., $\mathbf{X}$). The complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$ is denoted by $\mathcal{CN}(\mu,\sigma^{2})$. $\mathbf{I}$ is the identity matrix, $\odot$ denotes the Hadamard (element-wise) product, and * denotes convolution. Transpose and Hermitian transpose are represented by T and H, respectively. Flattened tensors are indicated with an overline (e.g., $\overline{\mathbf{X}}$). Attention scores are denoted as $\mathbf{A}_{(\cdot)}$, and extracted features as $\mathbf{F}_{(\cdot)}$ of a given layer as a subindex.
I System Model and Background
We consider a SIMO OFDM uplink communication system with $N_{BS}$ antennas at the Base Station (BS) as depicted in Fig. 1. A 3GPP Clustered Delay Line (CDL) channel model is applied in the frequency domain. Rather than just expressing received OFDM subcarriers, given that the transmitted resource grid (RG) is $\mathbf{X}_{\text{OFDM}}\in\mathbb{C}^{N_{\text{sym}}\times N_{\text{sc}}}$ where $N_{\text{sym}}$ is the number of symbols and $N_{\text{sc}}$ is the number of subcarriers in the OFDM system, we denote received RGs at the receiver of the BS as
$\mathbf{Y}_{\text{OFDM}}=\mathbf{H}\odot\mathbf{X}_{\text{OFDM}}+\mathbf{N},$ (1)
where $\mathbf{Y}_{\text{OFDM}}\in\mathbb{C}^{N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{BS}}}$ is the received RG at the BS, containing received symbols across subcarriers, OFDM symbols, and receive antennas. $\mathbf{H}\in\mathbb{C}^{N_{\text{sym }}\times N_{\text{sc }}\times N_{\text{BS}}}$ is the uplink SIMO channel in the frequency domain. Each element $\mathbf{H}_{k,s}^{(n)}$ represents the channel gain for the $k$-th symbol and $s$-th OFDM subcarrier at the $n$-th BS antenna. Finally, $\mathbf{N}\sim\mathcal{CN}\left(\mu,\sigma^{2}\mathbf{I}\right)$ is the AWGN noise at the BS.
From that point, after receiving OFDM RG, the goal of this study is to estimate the LLR values of received constellation points with DL techniques. However, before proceeding with DL model, we will give traditional receiver blocks required to estimate LLR values. The UE transmits pilot symbols at specific subcarriers and time slots, forming a pilot pattern in the OFDM RG $\mathbf{X}_{p}$. The received pilot signals at the BS are
$\mathbf{Y}_{p}=\mathbf{H}_{p}\odot\mathbf{X}_{p}+\mathbf{N}_{p},$ (2)
and the least squares (LS) channel estimate at the pilot locations is given by,
$\hat{\mathbf{H}}_{p}=\frac{\mathbf{Y}_{p}}{\mathbf{X}_{p}}$ (3)
Since pilots are available only at specific time-frequency positions, the full channel matrix $\hat{\mathbf{H}}$ is obtained by interpolation over the entire OFDM RG using linear interpolation. Following channel estimation, the BS performs linear minimum mean square error (LMMSE) equalization to mitigate fading and noise. The LMMSE equalization matrix for each symbol $k$ and subcarrier $s$ is
$\mathbf{W}_{k,s}^{\mathrm{LMMSE}}=\left(\hat{\mathbf{H}}_{k,s}^{\mathrm{H}}\hat{\mathbf{H}}_{k,s}+\frac{\sigma^{2}}{E_{s}}I\right)^{-1}\hat{\mathbf{H}}_{k,s}^{\mathrm{H}},$ (4)
and the equalized RG can be written as
$\hat{\mathbf{X}}_{\text{OFDM}}=\mathbf{W}^{\mathrm{LMMSE}}\mathbf{Y}_{\text{OFDM}}$ (5)
where $\mathbf{W}^{\mathrm{LMMSE}}\in\mathbb{C}^{N_{\text{sym }}\times N_{\text{sc }}\times N_{\mathrm{BS}}}$ is the equalization weight matrix for the entire RG and $\hat{\mathbf{X}}_{\text{OFDM}}\in\mathbb{C}^{N_{\text{sym }}\times N_{\text{sc}}}$ is the equalized symbol grid. After equalization, the equalized OFDM symbols are converted into bit-level LLRs. The LLR for the $i$-th bit considering a quadrature amplitude modulation (QAM) symbol is computed as:
$\mathbf{L}_{i}(k,s)=\log\frac{P\left(b_{i}=1\mid\hat{\mathbf{X}}^{k,s}_{\text{OFDM}}\right)}{P\left(b_{i}=0\mid\hat{\mathbf{{X}}}^{k,s}_{\text{OFDM}}\right)}$ (6)
where $\mathbf{L}\in\mathbb{R}^{N_{\text{sym }}\times N_{\text{sc }}\times N_{\text{bits }}}$ is the LLR grid, $P(b_{i}=1\mid\hat{\mathbf{X}}_{\text{OFDM}})$ and $P(b_{i}=0\mid\hat{\mathbf{X}}_{\text{OFDM}})$ are conditional probability functions for each possible bit values. Finally, the LLRs are fed to the low-density parity-check (LDPC) decoder, which decodes data and reconstructs the transmitted bits.
Figure 2: The architectures of proposed LLR estimation methods for neural receiver.
II Proposed DL-Based LLR Estimation Methods
In this section, we introduce two neural network architectures designed to directly estimate soft LLR values from the received OFDM RG $\mathbf{Y}$. The goal is to train a deep learning model that enhances BER-BLER performance by learning an effective mapping from the network input $\mathbf{Y_{\text{in}}}$ to the corresponding soft LLRs $\hat{\mathbf{L}}$. The details of this mapping function will be presented in the subsequent sections.
II-A Input Preprocessing and Embedding
The received complex-valued OFDM RG $\mathbf{Y}_{\text{OFDM}}\in\mathbb{C}^{N_{\text{sub}}\times N_{\text{sym}}\times N_{\mathrm{BS}}},$ is separated into real and imaginary components, producing a real-valued tensor, $\mathbf{Y}^{real}_{\text{OFDM}},\;\mathbf{Y}^{imag}_{\text{OFDM}}\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sym}}\times N_{\mathrm{BS}}}.$ In addition to the received signal samples, noise information in dB is appended as an auxiliary input, enabling the network to leverage the signal-to-noise ratio (SNR) context when producing LLRs. The noise information is broadcasted across the $(N_{\text{sym}},N_{\text{sc}},N_{\text{BS}})$ grid so that each network input token sees the same noise level value. Hence, the overall real-valued input $\mathbf{Y_{in}}$ can be denoted as
$\mathbf{Y_{in}}=[\mathbf{Y}_{\text{OFDM}}^{real},\mathbf{Y}_{\text{OFDM}}^{imag},\operatorname{log}_{10}(\boldsymbol{\mathrm{N_{0}}})],$ (7)
and has a shape of $\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times(2N_{\text{BS}}+1)}$.
II-B Dual Attention Transformer
The first architecture is a Dual Attention Transformer (DAT) designed to exploit both spatial (time-frequency) and channel dependencies (correlation between receive antennas) in the OFDM RG. Compared to a standard transformer encoder, DAT incorporates two parallel attention paths, one focusing on the spatial dimension of tokens and another focusing on the channel dimension to improve LLR estimation accuracy. Firstly, flattening the time–frequency axes of $\mathbf{Y_{in}}$ results,
$\overline{\mathbf{Y}}_{\text{in}}\in\mathbb{R}^{N_{T}\times N_{F}},\ N_{T}=N_{\text{sym}}N_{\text{sc}},\ N_{F}=2N_{\text{BS}}+1.$ (8)
A learned fully connected linear layer embeds these features into a D-dimensional model space,
$\mathbf{Z}=\text{FC}(\overline{\mathbf{Y}}_{\text{in}})=\overline{\mathbf{Y}}_{\text{in}}\mathbf{W}+\mathbf{b}\quad\in\mathbb{R}^{N_{T}\times N_{D}}$ (9)
where $\mathbf{W}$ denotes weights and $\mathbf{b}$ denotes bias and these are learnable parameters. Each block comprises Multi-Head Self-Attention (MHSA) and a Dual Attention Mechanism (DAM). Once tokenized, the data is fed through multiple Dual Attention Transformer Encoder (DAE) blocks.
II-B1 Multi-Head Self Attention
First of all, MHSA captures pairwise correlations among tokens, producing attention-weighted features from $\mathbf{Z}$,
$\mathbf{Z}_{(l)}=\operatorname{MHSA}(\mathbf{Z}_{(l-1)}).$ (10)
For an input sequence $\mathbf{Z}$, the self-attention mechanism projects $\mathbf{Z}$ using a learnable projection matrix $\mathbf{U}_{qkv}\in\mathbb{R}^{N_{T}\times 3N_{D}}$ into
$\mathbf{U}_{\mathrm{QKV}}$ $=\bigl{[}\mathbf{W}^{Q},\;\mathbf{W}^{K},\;\mathbf{W}^{V}\bigr{]},$ (11)
$[\mathbf{Q},\,\mathbf{K},\,\mathbf{V}]$ $=\mathbf{Z}\,\mathbf{U}_{\mathrm{QKV}}\;\in\;\mathbb{R}^{T\times 3D}.$
The resulting matrix separated into query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$ matrices where they are used to calculate attention scores by taking the scaled-dot product between the queries and keys. The softmax function is then applied row-wise to normalize the attention scores, ensuring that each row forms a probability distribution over the input tokens, and the result is multiplied by the $\mathbf{V}$ [17] which are expressed as
$\operatorname{MHSA}$ $=[\operatorname{head}_{1},\ldots,\operatorname{head}_{h}]\mathbf{W}^{O},$ (12)
where $\operatorname{head}_{i}$ $=\underbrace{\operatorname{softmax}\left(\mathbf{Q}_{i}\mathbf{K}_{i}^{T}/\sqrt{d_{k}}\right)\mathbf{V}_{i}}_{\text{scaled-dot product attention}},$
where $\mathbf{W}^{O}$ are trainable parameters of the heads. Finally, a residual connection and layer-norm applied before the DAM,
$\mathbf{U}=\operatorname{LN}(\mathbf{Z}_{l}+\mathbf{Z}_{(l-1)})\in\mathbb{R}^{N_{T}\times N_{D}}.$ (13)
II-B2 Dual-Attention Mechanism
DAM applies first a dense layer with number of D layers, and two parallel spatial (token-wise) and channel (feature-wise) attention operations. The computation of channel-wise attention given as
$\mathbf{A}_{c}=\operatorname{softmax}(\mathbf{W}_{c}\mathbf{U}^{\mathrm{T}}+\mathbf{b}_{c})\in\mathbb{R}^{N_{D}\times N_{D}},$ (14)
$\mathbf{V}_{c}=\mathbf{A}_{c}\mathbf{U}^{\mathrm{T}},\quad\mathbf{C}=\mathbf{V}_{c}^{\mathrm{T}}\in\mathbb{R}^{N_{T}\times N_{D}}.$ (15)
The computation of spatial-wise attention with T-units dense layer is given as
$\mathbf{A}_{s}=\operatorname{softmax}(\mathbf{U}\mathbf{W}_{s}+\mathbf{b}_{s})\in\mathbb{R}^{N_{T}\times N_{T}},$ (16)
$\mathbf{S}=\mathbf{A}_{s}\mathbf{U},\quad$ (17)
where $\mathbf{S}\in\mathbb{R}^{N_{T}\times N_{D}}$ and a layer normalization is added to the end of the DAM and it can be represented as
$\operatorname{DAM}(\mathbf{U})=$ $\operatorname{LN}(\mathbf{U}+\operatorname{\mathbf{A}}_{\operatorname{DAM}})=\mathbf{V}_{\operatorname{DAM}},$ (18)
$\text{where}\ \mathbf{A}_{\operatorname{DAM}}=$ $\mathbf{C}+{\mathbf{S}}.$
In the last part of the DAE, feed-forward networks (FFNs) are used to match the desired shape of the following DAE. These last FFNs also give flexibility to networks to plug in different shape parameters to play with networks to observe performance. Furthermore, to ensure stable gradients and to preserve both global and local level extracted features, skip connections are used in every internal block of the DAE; otherwise, the gradients would vanish due to the several attention mechanisms. Then, a $\operatorname{LN}$ is used to prevent internal covariance shift.
$\mathbf{Z_{out}}=\operatorname{reshape(}\operatorname{FC}(\operatorname{LN}(\operatorname{FFN}(\operatorname{PReLU}(\operatorname{FFN}(\mathbf{V}_{\operatorname{DAM}}))))))$ (19)
Finally, the last dense layer unit size is set to $N_{\text{bits}}$ in the last FC layer to match the desired shape $\mathbf{L}$, where the output is reshaped into $\mathbf{Z_{out}}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}$. In short, DAT emphasizes learning token-wise (spatial) and feature-wise (channel) correlations via self-attention and dual attention blocks. It is particularly adept at capturing long-range dependencies across subcarriers and symbols.
II-C Residual Dual Non-Local Attention Network
RDNLA combines the benefits of residual connections and non-local attention. Residual blocks capture local and multi-level features, while non-local blocks provide global dependency modeling. A dual strategy further splits attention into spatial-wise and channel-wise branches. Firstly, a convolutional layer extracts low-level features and maps the last channel into $C$ channel features as given,
$\mathbf{F_{0}}=\operatorname{PReLU}\bigl{(}\mathbf{W_{0}}*\mathbf{Y_{in}}+\mathbf{b}_{0}\bigr{)}\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C},$ (20)
where $\mathbf{Y_{in}}$ the input, as given in Eq. (7), $\mathbf{F_{0}}$ is the output of the layer. Also, $C$ is the growth rate parameter for the network which can be tuned. For easy interoperability, let’s denote this operation in Eq.(20) as $\operatorname{Conv_{1}}$ where the subindex is the number of convolutional layers applied.
II-C1 Parallel Residual Block
The network then applies a Parallel Residual Block (PRB), which merges features from two parallel arms, each containing Residual Dense Blocks (RDBs). Each RDB is defined as
$\mathbf{F}_{\text{RDB}}=\mathrm{Dropout}\bigl{(}\operatorname{\mathbf{F_{0}}+Conv_{2}}(\mathbf{F_{0}})\bigr{)}.$ (21)
Each arm in the PRB is composed of two sequential RDBs with skip connections around them. The outputs of the two arms are summed and then fused via a convolution, given as
$\mathbf{A}_{i}=\operatorname{RDB}^{(2)}(\operatorname{RDB}^{(1)}(\mathbf{F_{0}}))+\mathbf{F_{0}},\quad i=1,2,$ (22)
$\mathbf{F}_{\text{PRB}}=\operatorname{Conv}_{1}\bigl{(}\mathbf{A}_{1}+\mathbf{A}_{2}\bigr{)}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}.$ (23)
II-C2 Dual Non-Local Attention Block
Given an input feature tensor $\mathbf{X}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}$, Dual Non-Local Attention Block (DNLA) first applies three parallel $1\times 1$ convolutions to produce the lower-dimensional embeddings,
$\mathbf{\Theta}=\operatorname{Conv_{1}}(\mathbf{X}),\;\mathbf{\Phi}=\operatorname{Conv_{1}}(\mathbf{X}),\;\mathbf{G}=\operatorname{Conv_{1}}(\mathbf{X}),$ (24)
where their shape is $\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C^{\prime}}$ with $C^{\prime}=C/2$. By flattening the spatial dimensions into $N_{T}={N_{\text{sym}}\cdot N_{\text{sc}}}$ tokens, each of these feature maps is reshaped into a matrix of size $\mathbb{R}^{N_{T}\times C^{\prime}}$. In the spatial attention branch, we compute the attention scores as given,
$\mathbf{A}_{s}=\mathrm{softmax}\!\bigl{(}\overline{\mathbf{\Theta}}\overline{\mathbf{\Phi}}^{\mathrm{T}}\bigr{)}\;\in\;\mathbb{R}^{N_{T}\times N_{T}},$ (25)
where $\overline{\mathbf{\Theta}},\overline{\mathbf{\Phi}}$ is the flattened versions of embeddings and information is aggregated from all spatial locations,
$\mathbf{Z}_{s}=\mathbf{A}_{s}\overline{\mathbf{G}}\in\mathbb{R}^{N_{T}\times C^{\prime}}$ (26)
yielding a token matrix that is then reshaped back to $N_{\text{sym}}\times N_{\text{sc}}\times C^{\prime}$ and passed through a final $1\times 1$ convolution to restore the channel dimension to $C$.
$\mathcal{S}(\mathbf{X})=\operatorname{Conv_{1}}(\overline{\mathbf{Z}}_{s})\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}.$ (27)
Concurrently, we form the channel-wise attention scores as,
$\mathbf{A}_{c}=\mathrm{softmax}\!\bigl{(}\overline{\mathbf{\Theta}}^{\mathrm{T}}\,\overline{\mathbf{\Phi}}\bigr{)}\;\in\;\mathbb{R}^{C^{\prime}\times C^{\prime}},$ (28)
which are then aggregated to capture global channel correlations via
$\mathbf{Z}_{c}=\mathbf{A}_{c}\overline{\mathbf{G}}\in\mathbb{R}^{C^{\prime}\times N_{T}}.$ (29)
Later, the resulting output is projected back by
$\mathcal{C}(\mathbf{X})=\operatorname{Conv_{1}}(\operatorname{reshape}(\mathbf{Z}_{c}^{\mathrm{T}}))\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}.$ (30)
Lastly, the spatial and channel attention paths are summed and added to the original input via a residual connection,
$\mathbf{F}_{\mathrm{DNLA}}=\mathbf{X}\;+\;\mathcal{S}(\mathbf{X})\;+\;\mathcal{C}(\mathbf{X}),\;\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C},$ (31)
where $\mathcal{S}(\mathbf{X})$ and $\mathcal{C}(\mathbf{X})$ denote the restored-channel outputs of the spatial and channel branches, respectively. This dual-path design allows each location to attend to all others across both spatial and channel dimensions, while the skip connection preserves identity information. The network’s output can be obtained by adding the output of the first convolutional layer to the output of each block and then passing through the last convolutional layer, as shown below:
$\mathbf{Z}_{\text{out }}=\operatorname{Conv_{1}}\left(\mathbf{F}_{\mathrm{PRB}_{2}}+\mathbf{F}_{0}\right)\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}.$ (32)
In summary, DAT and RDNLA predict the soft-LLR values for given batch examples,
$\hat{\mathbf{L}}=f_{\theta}(\mathbf{Y_{in}})\in\mathbb{R}^{N_{\text{batch}}\times N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{bits}}},$ (33)
and each element of $\hat{\mathbf{L}}$ corresponds to the LLR of the $i$-th coded bit at symbol $k$, subcarrier $s$, and batch sample $m$.
II-D Model Training
The given models are trained using the bit-metric decoding (BMD) rate, which serves as an information theoretic objective for bit-interleaved coded modulation (BCIM) systems. The BMD rate is computed from the transmitted coded bits and the predicted LLRs, and its optimization ensures accurate soft-bit estimation for subsequent decoding. Binary-cross entropy (BCE) loss function compares the predicted LLRs with the true transmitted coded bits over the entire OFDM RG, which is given by
$\mathcal{L}_{\mathrm{BCE}}(B,\hat{L})=\mathbb{E}[B\log\sigma(\hat{L})+(1-B)\log(1-\sigma(\hat{L}))],$ (34)
where $B$ represents the transmitted ground truth bits and $\hat{L}$ represents the predicted LLRs, finally $\sigma(\hat{L})$ is the sigmoid activation ensuring LLR interpretation as probabilities. We used the expectation operator $\mathbb{E}[\cdot]$ in loss function for conciseness, which expresses averaging over all dimensions. We used the Adam optimizer for gradients. The goal of the model is to learn the optimal parameters that minimize the loss function in Eq. (34), formulated as
$\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\mathrm{BCE}}\left(B,f_{\theta}\left(\mathbf{Y_{in}}\right)\right),$ (35)
where $\theta$ represents the learnable parameters of the model and $f_{\theta}(\cdot)$ is the learned mapping function, also known as the soft LLR estimator.
Figure 3: BER (a) and BLER (b) of the proposed neural receivers with state-of-the art models at varying noise levels.
III Simulation Experiments and Results
This section presents the simulation experiments and results using Nvidia’s open-source Sionna library. A wireless communication system consisting of a single antenna UE ($N_{\mathrm{UE}}=1$) and a BS with two antennas $N_{\mathrm{BS}}=2$ in Fig. 1 is generated within the Sionna simulator using a single stream configuration. The OFDM system consists of 76 subcarriers and 14 OFDM symbols per each frame, where 2 symbols are allocated for pilots. A cyclic prefix of 6 samples is used to mitigate inter-symbol interference. The signal is modulated using QAM, while LDPC encoding and decoding are applied at a coding rate of $R_{C}=0.5$. The proposed two neural receiver architectures, namely DAT and RDNLA, are implemented within the Sionna simulator in addition to eight existing SOTA DL architectures from the literature. BER and BLER results for ten architectures are shown as $E_{b}/N_{0}$ is swept from -3 dB to 5 dB during simulations, however the plots confined to specific interval for clarity. Each performance result corresponds to the average of $3200$ Monte-Carlo experiments, with the target block error run set to 1000. For training, the batch size is set to 128 while the number of epochs is $10^{5}$.
Fig. 3(a) shows the BER results of ten different receiver models including our proposed DAT and RDNLA models under different $E_{b}/N_{0}$ values. The figure distinguishes two groups of receiver architectures. The first group (ResNet [6], Transformer [16], ResDNLA, DAT) achieves the BER results below $10^{-4}$ even $E_{b}/N_{0}$ is lower than 0.5 dB. This success is due to their architectural components such as local, global attention blocks, and residual connections, which effectively capture complex and global relationships and enhance feature representation for neural receiver tasks. The second group (EffNetV2, NAFNet ResUNet, RIDNet[18, 19, 20, 21]) have demonstrated lower BER compared to their counterpart, especially EffNet-V2, NAFNet and ResUNet. Although these models perform well in semantic segmentation and incorporate skip connections and conventional attention mechanisms, they underperform in LLR estimation due to their inability to effectively capture both local and global dependencies required for this task. The first group excels at modelling both time-frequency and spatial–channel relationships, while DAT and RDNLA further outperform standard Transformer and ResNet architectures by using finely crafted attention mechanisms that effectively capture complex feature dependencies and preserve global features. DAT and RDNLA are able to achieve $10^{-4}$ BER below 0 dB while Transformer and Resnet can achieve this result above this point. This can be seen as a minor gain however trend shows that in higher SNR regimes DAT and RDNLA draw away especially after 0 dB because Transformer and ResNet are not able to learn local dependencies at these points.
Fig. 3(b) shows the BLER of the mentioned receiver models including our proposed DAT and RDNLA models under different $E_{b}/N_{0}$ values. The BLER results are distinguished from the BER results. The BLER values of Transformer and ResNet are struggling to converge at high SNR compared to their proposed counterparts DAT and RDNLA. This shows again Transformer and ResNet are more prone to block-level errors, due to their lack of specific attention mechanisms that result in errors on subsequent element prediction. The proposed models further exploit the complex features and can correct both global and subsequent local information.
TABLE I: Mean inference time of models per OFDM frame
Model Time (ms)
Transformer (Encoder) 11.699
NAFNet 11.891
ResNet 12.032
EffNet-V2 12.270
DAT 12.366
RDNLA 12.684
ResUNet 12.910
RIDNet 13.349
Table I shows that all models process an OFDM frame within a narrow window (±1.65 ms), reflecting a balanced design in layer depth and attention complexity. The Transformer is the fastest, as it relies on dense layers and self-attention, both dominated by efficient matrix multiplications. In contrast, convolutional models (ResNet, ResUNet, RIDNet, RDNLA) use many small kernel operations that are memory-bound.
IV Discussion and Future Work
This paper proposes two neural network–based OFDM receivers and compares their BER and BLER performance with eight existing state-of-art models, where the DAT employs a transformer architecture with an attention mechanism while the RDNLA utilizes a parallel residual architecture with a non-local attention block. The simulation experiments are conducted using the open-source Sionna simulator. The results show that the proposed models outperform both traditional receiver systems and existing neural receivers. Also, results show that BER alone is not enough to train the model because it ignores how errors cluster, a factor that heavily impacts BLER. Models that disperse bit errors tend to achieve lower BLER even with higher BER because scattered errors are easier for the decoder to correct, whereas clustered errors can overwhelm it and trigger entire block failures. Thus, we will analyze the spatial and temporal patterns of decoding errors and design architectures and training schemes that promote error dispersion. In particular, we plan to incorporate error-aware loss functions and decoder-in-the-loop feedback to directly optimize for BLER.

### v3

@@
-Deep Neural OFDM Receivers: Two Novel Architectures for BER-BLER Optimization and Comparison with State-of-the-Art Architectures
+Novel Deep Neural OFDM Receiver Architectures for LLR Estimation
 Abstract
@@
 The simulation results show that DAT and RDNLA outperform both traditional communication systems and existing neural receiver models.
-The computational efficiency of the proposed neural receivers supports their feasibility for next-generation communication systems.
 Index Terms:
 Neural Receivers, LLR, Transformers
+I Introduction
 Data-driven techniques such as machine learning (ML) and deep learning (DL) have seen significant advances recently, not only in foundational algorithms but also in hardware design perspective, mainly driven by the “artificial intelligence (AI) at the edge” concept.
 For example, tensor accelerators along with new AI algorithms make computationally challenging tasks feasible to be deployed in real-world applications.
-These advancements facilitate an increasing number of AI deployements at the edge [1].
-Wireless communication community has also been heavily investing AI/ML based applications, where they can either build new foundational models or adopt existing ones from other domains by training them using wireless communication datasets.
-Considering advanced radio access network (RAN) technologies along with the usage of larger bandwidths in the upcoming 3GPP releases, signal processing requirements of enormous data in real-time make a traditional mathematical-based system design with cascaded communication blocks consisting heuristic algorithms prohibitively expensive.
+These advancements facilitate an increasing number of AI deployments at the edge [1].
+The wireless communication community has also been heavily investing in AI/ML-based applications, where they can either build new foundational models or adapt existing ones from other domains by training them using wireless communication datasets.
+Considering advanced radio access network (RAN) technologies along with the usage of larger bandwidths in the upcoming 3GPP releases, signal processing requirements of enormous data in real-time make a traditional mathematical-based system design with cascaded communication blocks consisting of heuristic algorithms prohibitively expensive.
 This brings an AI-native system design forefront, where the AI/ML approaches can focus on the most important features of the underlying system by appropriately setting the weights of the trained models which provide a direct mapping between transmitted (input) and received (output) information.
@@
 Furthermore, it is shown in [6] that an end-to-end neural receiver for OFDM systems under realistic fading channels can significantly reduce pilot overhead by providing a similar bit-error rate (BER) performance compared to conventional pilot-based systems.
-The recent studies on a tensor native wireless communication [7] and ray tracing based simulation framework Sionna [8] provide important advancements in the context of neural receivers.
+The recent studies on a tensor native wireless communication [7] and Nvidia’s ray tracing based simulation framework Sionna [8] provide important advancements in the context of neural receivers.
 The potential of neural receiver architectures has been demonstrated in [9] and [10], where neural receivers are utilized in 5G multi-user multiple input-multiple output (MIMO) scenarios in real-time while adhering to the standard compliance.
@@
 Proposed models also demonstrate notable improvements over other state-of-the-art (SOTA) models.
+Additionally, we provided the computation times of each model with the conventional least square (LS) estimation receiver.
   • We then explore the architecture of these models, carefully going over each element of their construction.
 We highlight how each element contributes to feature extraction and learning at the global and local levels that eventually affect BER and BLER.
-  • Lastly, we will make publicly available the source code of our proposed models together with other SOTA model implementations [^1] https://github.com/Hisar-Research/sionna-nn-receiver-playground.
+  • Lastly, we will make the source code of our proposed models together with other SOTA model implementations publicly available [^1] https://github.com/Hisar-Research/sionna-nn-receiver-playground.
 This allows researchers and practitioners to explore, test, and expand upon our findings.
@@
 Attention scores are denoted as $\mathbf{A}_{(\cdot)}$, and extracted features as $\mathbf{F}_{(\cdot)}$ of a given layer as a subindex.
-I System Model and Background
+II System Model and Background
 We consider a SIMO OFDM uplink communication system with $N_{BS}$ antennas at the Base Station (BS) as depicted in Fig.
@@
 From that point, after receiving OFDM RG, the goal of this study is to estimate the LLR values of received constellation points with DL techniques.
-However, before proceeding with DL model, we will give traditional receiver blocks required to estimate LLR values.
+However, before proceeding with the DL model, we will give traditional receiver blocks required to estimate LLR values.
 The UE transmits pilot symbols at specific subcarriers and time slots, forming a pilot pattern in the OFDM RG $\mathbf{X}_{p}$.
@@
 $\mathbf{Y}_{p}=\mathbf{H}_{p}\odot\mathbf{X}_{p}+\mathbf{N}_{p},$ (2)
-and the least squares (LS) channel estimate at the pilot locations is given by,
+and the LS channel estimate at the pilot locations is given by,
 $\hat{\mathbf{H}}_{p}=\frac{\mathbf{Y}_{p}}{\mathbf{X}_{p}}$ (3)
@@
 Finally, the LLRs are fed to the low-density parity-check (LDPC) decoder, which decodes data and reconstructs the transmitted bits.
-Figure 2: The architectures of proposed LLR estimation methods for neural receiver.
-II Proposed DL-Based LLR Estimation Methods
+Figure 2: The DAT (left) and RDNLA (right) LLR estimation architectures for neural receiver.
+III Proposed DL-Based LLR Estimation Methods
 In this section, we introduce two neural network architectures designed to directly estimate soft LLR values from the received OFDM RG $\mathbf{Y}$.
@@
 The details of this mapping function will be presented in the subsequent sections.
-II-A Input Preprocessing and Embedding
+III-A Input Preprocessing and Embedding
 The received complex-valued OFDM RG $\mathbf{Y}_{\text{OFDM}}\in\mathbb{C}^{N_{\text{sub}}\times N_{\text{sym}}\times N_{\mathrm{BS}}},$ is separated into real and imaginary components, producing a real-valued tensor, $\mathbf{Y}^{real}_{\text{OFDM}},\;\mathbf{Y}^{imag}_{\text{OFDM}}\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sym}}\times N_{\mathrm{BS}}}.$ In addition to the received signal samples, noise information in dB is appended as an auxiliary input, enabling the network to leverage the signal-to-noise ratio (SNR) context when producing LLRs.
@@
 and has a shape of $\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times(2N_{\text{BS}}+1)}$.
-II-B Dual Attention Transformer
+III-B Dual Attention Transformer
 The first architecture is a Dual Attention Transformer (DAT) designed to exploit both spatial (time-frequency) and channel dependencies (correlation between receive antennas) in the OFDM RG.
@@
 Once tokenized, the data is fed through multiple Dual Attention Transformer Encoder (DAE) blocks.
-II-B1 Multi-Head Self Attention
+III-B1 Multi-Head Self Attention
 First of all, MHSA captures pairwise correlations among tokens, producing attention-weighted features from $\mathbf{Z}$,
@@
 $\mathbf{U}_{\mathrm{QKV}}$ $=\bigl{[}\mathbf{W}^{Q},\;\mathbf{W}^{K},\;\mathbf{W}^{V}\bigr{]},$ (11)
-$[\mathbf{Q},\,\mathbf{K},\,\mathbf{V}]$ $=\mathbf{Z}\,\mathbf{U}_{\mathrm{QKV}}\;\in\;\mathbb{R}^{T\times 3D}.$
+$[\mathbf{Q},\,\mathbf{K},\,\mathbf{V}]$ $=\mathbf{Z}\,\mathbf{U}_{\mathrm{QKV}}\;\in\;\mathbb{R}^{N_{T}\times 3N_{D}}.$
 The resulting matrix separated into query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$ matrices where they are used to calculate attention scores by taking the scaled-dot product between the queries and keys.
@@
 $\mathbf{U}=\operatorname{LN}(\mathbf{Z}_{l}+\mathbf{Z}_{(l-1)})\in\mathbb{R}^{N_{T}\times N_{D}}.$ (13)
-II-B2 Dual-Attention Mechanism
+III-B2 Dual-Attention Mechanism
 DAM applies first a dense layer with number of D layers, and two parallel spatial (token-wise) and channel (feature-wise) attention operations. The computation of channel-wise attention given as
@@
 $\mathbf{Z_{out}}=\operatorname{reshape(}\operatorname{FC}(\operatorname{LN}(\operatorname{FFN}(\operatorname{PReLU}(\operatorname{FFN}(\mathbf{V}_{\operatorname{DAM}}))))))$ (19)
-Finally, the last dense layer unit size is set to $N_{\text{bits}}$ in the last FC layer to match the desired shape $\mathbf{L}$, where the output is reshaped into $\mathbf{Z_{out}}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}$.
+Finally, the last dense layer unit size is set to $N_{\text{bits}}$ in the last FC layer to match the desired shape $\mathbf{L}$, where the output is reshaped into $\mathbf{Z_{out}}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{bits}}}$.
 In short, DAT emphasizes learning token-wise (spatial) and feature-wise (channel) correlations via self-attention and dual attention blocks.
 It is particularly adept at capturing long-range dependencies across subcarriers and symbols.
-II-C Residual Dual Non-Local Attention Network
+III-C Residual Dual Non-Local Attention Network
 RDNLA combines the benefits of residual connections and non-local attention.
@@
 Also, $C$ is the growth rate parameter for the network which can be tuned.
-For easy interoperability, let’s denote this operation in Eq.(20) as $\operatorname{Conv_{1}}$ where the subindex is the number of convolutional layers applied.
-II-C1 Parallel Residual Block
+For easy interoperability, let’s denote this operation in Eq.
+(20) as $\operatorname{Conv_{1}}$ where the subindex is the number of convolutional layers applied.
+III-C1 Parallel Residual Block
 The network then applies a Parallel Residual Block (PRB), which merges features from two parallel arms, each containing Residual Dense Blocks (RDBs). Each RDB is defined as
@@
 $\mathbf{F}_{\text{PRB}}=\operatorname{Conv}_{1}\bigl{(}\mathbf{A}_{1}+\mathbf{A}_{2}\bigr{)}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}.$ (23)
-II-C2 Dual Non-Local Attention Block
+III-C2 Dual Non-Local Attention Block
 Given an input feature tensor $\mathbf{X}\in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times C}$, Dual Non-Local Attention Block (DNLA) first applies three parallel $1\times 1$ convolutions to produce the lower-dimensional embeddings,
@@
 This dual-path design allows each location to attend to all others across both spatial and channel dimensions, while the skip connection preserves identity information.
-The network’s output can be obtained by adding the output of the first convolutional layer to the output of each block and then passing through the last convolutional layer, as shown below:
-$\mathbf{Z}_{\text{out }}=\operatorname{Conv_{1}}\left(\mathbf{F}_{\mathrm{PRB}_{2}}+\mathbf{F}_{0}\right)\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times 2}.$ (32)
+The network’s output can be obtained by adding the output of the first convolutional layer to the output of each block and then passing through the last convolutional layer with $N_{\text{bits}}$ filters, as shown below
+$\mathbf{Z}_{\text{out }}=\operatorname{Conv_{1}}\left(\mathbf{F}_{\mathrm{PRB}_{2}}+\mathbf{F}_{0}\right)\ \in\mathbb{R}^{N_{\text{sym}}\times N_{\text{sc}}\times N_{\text{bits}}}.$ (32)
 In summary, DAT and RDNLA predict the soft-LLR values for given batch examples,
@@
 and each element of $\hat{\mathbf{L}}$ corresponds to the LLR of the $i$-th coded bit at symbol $k$, subcarrier $s$, and batch sample $m$.
-II-D Model Training
+III-D Model Training
 The given models are trained using the bit-metric decoding (BMD) rate, which serves as an information theoretic objective for bit-interleaved coded modulation (BCIM) systems.
 The BMD rate is computed from the transmitted coded bits and the predicted LLRs, and its optimization ensures accurate soft-bit estimation for subsequent decoding.
-Binary-cross entropy (BCE) loss function compares the predicted LLRs with the true transmitted coded bits over the entire OFDM RG, which is given by
+The binary-cross entropy (BCE) loss function compares the predicted LLRs with the true transmitted coded bits over the entire OFDM RG, which is given by
 $\mathcal{L}_{\mathrm{BCE}}(B,\hat{L})=\mathbb{E}[B\log\sigma(\hat{L})+(1-B)\log(1-\sigma(\hat{L}))],$ (34)
@@
 where $\theta$ represents the learnable parameters of the model and $f_{\theta}(\cdot)$ is the learned mapping function, also known as the soft LLR estimator.
-Figure 3: BER (a) and BLER (b) of the proposed neural receivers with state-of-the art models at varying noise levels.
-III Simulation Experiments and Results
+Figure 3: BER (a) and BLER (b) of the proposed neural receivers with SOTA models at varying noise levels.
+IV Simulation Experiments and Results
 This section presents the simulation experiments and results using Nvidia’s open-source Sionna library.
 A wireless communication system consisting of a single antenna UE ($N_{\mathrm{UE}}=1$) and a BS with two antennas $N_{\mathrm{BS}}=2$ in Fig.
-1 is generated within the Sionna simulator using a single stream configuration.
-The OFDM system consists of 76 subcarriers and 14 OFDM symbols per each frame, where 2 symbols are allocated for pilots.
+1 is generated using a single stream configuration.
+The OFDM system consists of 76 subcarriers and 14 OFDM symbols per frame, where 2 symbols are allocated for pilots.
 A cyclic prefix of 6 samples is used to mitigate inter-symbol interference.
 The signal is modulated using QAM, while LDPC encoding and decoding are applied at a coding rate of $R_{C}=0.5$.
-The proposed two neural receiver architectures, namely DAT and RDNLA, are implemented within the Sionna simulator in addition to eight existing SOTA DL architectures from the literature.
-BER and BLER results for ten architectures are shown as $E_{b}/N_{0}$ is swept from -3 dB to 5 dB during simulations, however the plots confined to specific interval for clarity.
+The proposed two neural receiver architectures, namely DAT and RDNLA, are implemented in addition to eight existing SOTA DL architectures from the literature.
+BER and BLER for all ten architectures are plotted as functions of $E_{b}/N_{0}$; for clarity, the BER curves span from -2.5 dB to 1.5 dB, while the BLER curves cover -1.5 dB to 2 dB.
 Each performance result corresponds to the average of $3200$ Monte-Carlo experiments, with the target block error run set to 1000.
@@
 3(a) shows the BER results of ten different receiver models including our proposed DAT and RDNLA models under different $E_{b}/N_{0}$ values.
-The figure distinguishes two groups of receiver architectures.
+Two separate performance clusters of receiver architectures emerge in the figure.
 The first group (ResNet [6], Transformer [16], ResDNLA, DAT) achieves the BER results below $10^{-4}$ even $E_{b}/N_{0}$ is lower than 0.5 dB.
-This success is due to their architectural components such as local, global attention blocks, and residual connections, which effectively capture complex and global relationships and enhance feature representation for neural receiver tasks.
-The second group (EffNetV2, NAFNet ResUNet, RIDNet[18, 19, 20, 21]) have demonstrated lower BER compared to their counterpart, especially EffNet-V2, NAFNet and ResUNet.
+This success is due to their architectural components such as local, and global attention blocks, and residual connections, which effectively capture complex and global relationships and enhance feature representation for neural receiver tasks.
+The second group (EffNet-V2, NAFNet ResUNet, RIDNet [18, 19, 20, 21]) have demonstrated lower BER compared to their counterpart, especially EffNet-V2, NAFNet and ResUNet.
 Although these models perform well in semantic segmentation and incorporate skip connections and conventional attention mechanisms, they underperform in LLR estimation due to their inability to effectively capture both local and global dependencies required for this task.
 The first group excels at modelling both time-frequency and spatial–channel relationships, while DAT and RDNLA further outperform standard Transformer and ResNet architectures by using finely crafted attention mechanisms that effectively capture complex feature dependencies and preserve global features.
-DAT and RDNLA are able to achieve $10^{-4}$ BER below 0 dB while Transformer and Resnet can achieve this result above this point.
-This can be seen as a minor gain however trend shows that in higher SNR regimes DAT and RDNLA draw away especially after 0 dB because Transformer and ResNet are not able to learn local dependencies at these points.
+DAT and RDNLA achieve a BER of $10^{-4}$ at $E_{b}/N_{0}$ below 0 dB.
+In contrast, the Transformer and ResNet require $E_{b}/N_{0}$ above 0 dB to reach the same.
+While the gain is modest at low $E_{b}/N_{0}$ (below -1 dB), beyond 0 dB DAT and RDNLA steadily pull ahead, as the Transformer and ResNet fall into decline in capturing local dependencies at higher SNR.
 Fig.
@@
 The proposed models further exploit the complex features and can correct both global and subsequent local information.
-TABLE I: Mean inference time of models per OFDM frame
+TABLE I: Mean inference times per OFDM RG
 Model Time (ms)
@@
 RIDNet 13.349
-Table I shows that all models process an OFDM frame within a narrow window (±1.65 ms), reflecting a balanced design in layer depth and attention complexity.
+LS-Estimation 0.14
+Table I shows that all models process an OFDM RG within a narrow window (±1.65 ms), reflecting a balanced design in layer depth and attention complexity.
 The Transformer is the fastest, as it relies on dense layers and self-attention, both dominated by efficient matrix multiplications.
 In contrast, convolutional models (ResNet, ResUNet, RIDNet, RDNLA) use many small kernel operations that are memory-bound.
-IV Discussion and Future Work
+Not surprisingly, the conventional LS-estimation receiver is faster than the neural models, underscoring the trade-off between computational efficiency and performance.
+V Discussion and Future Work
 This paper proposes two neural network–based OFDM receivers and compares their BER and BLER performance with eight existing state-of-art models, where the DAT employs a transformer architecture with an attention mechanism while the RDNLA utilizes a parallel residual architecture with a non-local attention block.
-The simulation experiments are conducted using the open-source Sionna simulator.
 The results show that the proposed models outperform both traditional receiver systems and existing neural receivers.
-Also, results show that BER alone is not enough to train the model because it ignores how errors cluster, a factor that heavily impacts BLER.
-Models that disperse bit errors tend to achieve lower BLER even with higher BER because scattered errors are easier for the decoder to correct, whereas clustered errors can overwhelm it and trigger entire block failures.
+However, neural receivers still require efficient architectures to catch the speed of conventional receivers.
+In addition, the results show that BER alone is insufficient to train the model because it ignores how errors cluster, a factor that heavily impacts BLER.
+The models that disperse bit errors tend to achieve lower BLER even with higher BER because scattered errors are easier for the decoder to correct, whereas the clustered errors can overwhelm it and trigger entire block failures.
 Thus, we will analyze the spatial and temporal patterns of decoding errors and design architectures and training schemes that promote error dispersion.
-In particular, we plan to incorporate error-aware loss functions and decoder-in-the-loop feedback to directly optimize for BLER.
+In particular, we plan to incorporate error-aware loss functions and decoder-in-the-loop feedback to directly optimize BLER.